I have to put this in standard form....
3a+b-2=b

Thank you!

Answer by jim_thompson5910(35256)

3a+b-2=b

3a+b-2-b = 0

3a + 0b - 2 = 0

3a + 0b = 2

So if a and b are your variables, then standard form would be ka + mb = n where k, m, n are fixed numbers. So because 3a + 0b = 2 fits that form, 3a + 0b = 2 is in standard form.